Hi
Brian had a sum of money. He spent an equal amount of money each day. After 6 days he had 3/5 of his money left. 3 days later he had 360 dollars left. How much did he have at first.
Thanks

Answer by ikleyn(52908)

.
Brian had a sum of money. He spent an equal amount of money each day.
After 6 days he had 3/5 of his money left. 3 days later he had 360 dollars left.
How much did he have at first.
~~~~~~~~~~~~

From the condition, you see that Brian spent  1 - 3/5 = 2/5  of his money in 6 day.


Keeping the same rate, in the next 3 days he will spend half of 2/5, which is 1/5.


So,  in 6 + 3 = 9 days  Brian will spend  2/5 + 1/5 = 3/5  of his money.


Hence, after 9 days Brian will still have  1- 3/5 = 2/5  of his money left.


Thus 2/5 of his money is 360 dollars.


Hence,  1/5  of his money is half of  360  dollars,  i.e.  180 dollars.


It means that  Brian had initially  5 times as much as 180 dollars, i.e.  900  dollars.    ANSWER

Solved (mentally, by reasoning, without using equations).